A company’s board of directors wants to form a committee of 2 members. 8 members to choose from. How many different committees of 2 members could be formed?
Answer by rothauserc(4718)

We want the total number of combinations of 8 members taken 2 at a time
:
8C2 = 8!/(2! * (8-2)!) = (8 * 7)/2 = 28 different committees of 2 members
: